# Emergent predictability in microbial ecosystems

### v2

@@ II Results
Figure 2: Standard ecological models exhibit self-averaging, but not the nuanced forms of emergent simplicity defined here. As an example, we consider the problem of predicting ecosystem resistance to an invader. A, B: Applying our framework to simulated data from two standard ecological models: the Generalized Lotka-Volterra model (panel A) and the Consumer Resource model (panel B), with the observable of interest $Y_{\mu}$ defined as the post-invasion abundance of a designated “invader” strain. Shown are the Pareto fronts in the error-complexity plane, for three tiers of community richness. The Pareto fronts show no signature of either emergent coarse-grainability or emergent predictability sensu Fig. 1D. C: Same data as in B, replotted with the $y$ axis showing the absolute, rather than fractional, variance of the observable $Y$. In these axes, the Pareto fronts show a clear shift downwards. This corresponds to the self-averaging behavior (as community richness increases, the distribution of the observable $Y$ becomes more narrow). We conclude that, unlike the simple self-averaging behavior, both forms of emergent simplicity defined here are not a typical feature of standard ecological models.
II.1 Standard ecological models do not generically exhibit either form of emergent simplicity
Before we apply our framework to real data, it is important to establish whether observing the behaviors of Fig. 1D would be surprising [32]. For this, we consider two classic models widely used in theoretical ecology: the Generalized Lotka-Volterra model (GLV) with a random interaction matrix, and a Consumer-Resource model (CRM) with a random resource consumption matrix.
To apply our framework, we need to specify the property $Y$ we would seek to predict. Here, we focus on community invadeability by a specified strain. This choice is convenient because it is a relevant and non-trivial community property that can be accessed in both modeling frameworks.
The exact simulation protocol and parameter choices are described in the SI, but our overall approach is similar for both models, and proceeds as follows: (1) Construct a species pool, by starting from a random instance of a larger ecosystem and simulating its dynamics to obtain a coexisting set of about 25 species. (2) Define an “invader,” drawing a random new species from the same ensemble. (3) Generate the synthetic dataset. We simulate the experimental protocol of Fig. 1A, with 1000 samples for each of three tiers of community richness: low richness, 1-5 strains; middle richness, 10-15 strains; and high richness, 20-25 strains. These tiers were chosen to match the empirical dataset of Ref. [33] used below. For each community, we first record the equilibrium abundances of all strains as $n_{i\mu}$, and then simulate the outcome of its invasion by the pathogenic strain, recording the post-invasion abundance of the pathogen as $Y_{\mu}$. Finally, we (4) construct the information-prediction Pareto fronts defined above, using a heuristic algorithm to search the space of possible coarsenings (for details, see SI section .6). The results, averaged over 10 random choices of the invader, are presented in Fig. 2A (GLV model) and Fig. 2B (CRM model).
Figure 3: Both available empirical datasets (Refs. [33] and [34]) exhibit emergent predictability. A-E: The Pareto fronts in error-complexity plane, computed for five properties of interest. Lines show the fraction of variance left unexplained by the best model within a given tier of description complexity, when dataset is restricted to communities of low richness (orange), or high richness (purple). At high richness, the Pareto front shifts down, indicating that simpler descriptions become more predictive (emergent predictability, Fig. 1E). Each point and error bar shows the median quartiles across 100 random 50-50 splits of the data into training and testing sets. G,H: Illustrating the result by plotting predicted versus measured butyrate concentration (the points on the Pareto front indicated by the arrowheads in A). At high richness, a simple compositional feature explains 85% of variance in butyrate production (panel G). In contrast, for low-richness communities, even the most predictive model of comparable complexity explains only 49% of variance (panel F).
We make two observations. First, we see that models with $I(\Psi)\gtrsim 3$ bits are generally able to capture $\gtrsim 75\%$ of variation in these simulated data. (Note that we simulated 1000 samples, and as a result there is no overfitting signature: the Pareto fronts show no uptick at high complexity). Even in the absence of simulated experimental noise, this level of performance was not guaranteed. Thus, Fig. 2 confirms that the model class of linear regressions, while simple, can be sufficiently expressive to describe datasets structured by ecological interactions. The one notable exception is the high-diversity GLV, where linear regressions provide no explanatory power.
Second, we see that in both cases, the Pareto fronts shift up and right, not down (Fig. 1E) or left (Fig. 1D). This analysis confirms our intuition that the emergent simplicity as defined here is not a behavior one would generically expect. Contrast this to Fig. 2C, where the same CRM simulation data are replotted using the absolute prediction error rather than relative. Here, the Pareto fronts exhibit a clear shift downwards. This is the phenomenon of ‘emergent reproducibility’ or functional convergence, where the distribution of $Y$ gets narrower as richness increases. The contrast with Fig. 2A, B highlights that both forms of emergent simplicity shown in Fig. 1D, E are distinct from this more familiar scenario, widely reported both empirically [4, 5, 6, 21, 7, 10, 22, 11, 12, 13] and in models [8, 30]. In Fig. 1D, E and Fig. 2A, B, the prediction error of $Y$ is always normalized to the overall width of the distribution of $Y$ across examples of the same richness tier. Our goal is to quantify forms of emergent simplicity going beyond emergent reproducibility, and our normalization explicitly removes this simpler effect.
To summarize, Fig. 2 demonstrates that standard ecological models do not generically exhibit the nuanced forms of emergent simplicity defined here. Thus, unlike the simple self-averaging, observing either of these phenomena in empirical data would be surprising.
II.2 Empirical data exhibit emergent predictability
@@
To test the emergent simplicity hypothesis in data, we need a dataset satisfying several conditions. First, it must follow the combinatorial assembly protocol of Fig. 1A, with communities assembled from a fixed pool of strains. Surveying the literature, we identified 9 candidate datasets (across both microbial and non-microbial contexts; see Supp Table S1). However, for our purposes it is essential that the assayed communities span a range of richness, so that at least two tiers of community diversity can be defined and compared. Additionally, if $N$ is the number of samples and $S$ the number of strains in the pool, $N/S$ should be sufficiently large (for each richness tier) to allow training the model from our chosen model class. Our choice of linear regressions as model class makes this last condition the most permissive, and identified two datasets whose structure allows our question to be asked. These are the data of Clark et al. [33], who assayed communities of 1-25 gut-derived taxa and measured the production of butyrate, acetate, lactate and succinate; and the work of Kehe et al. [34], who used a droplet-based platform to screen communities of 1-19 isolates, measuring the abundance of a model plant symbiont H. frisingense as the property of interest. Together, these two datasets provide measurements of five properties we can examine for evidence of emergent simplicity.
As above, for each dataset, we resolve three tiers of community richness (with the high-richness tier, $>20$ strains, only available for Clark et al. data). Since Clark et al. did not directly measure absolute strain abundances in the community (only its total OD and the 16S makeup), we adopted an approach where the relative abundances estimated from the 16S data were first scaled by community OD and then standardized to zero mean and unit variance before training any regression models. For consistency, the same standardization procedure was also applied to Kehe et al data; though omitting this step does not qualitatively change the outcome in either case; see SI fig. S3. We then applied the same heuristic algorithm as above to infer the respective Pareto front. The results are shown in Fig. 3A-E.
Remarkably, we find that four out of five properties exhibit emergent predictability (but not emergent coarse-grainability): simple models become increasingly predictive in communities of higher richness. As a control, this behavior is not observed for measurements of community optical density (OD), or if the values of $Y$ or species abundances are randomly permuted (see SI), indicating that the behavior observed in Fig. 3 is not an artifact of our analysis.
The additional controls we performed include subsampling the data, perturbing the groupings forming the putative Pareto front, and considering alternative metrics of complexity $I(\Psi)$; see SI. For all four properties exhibiting emergent predictability in Fig. 3, the result is robust against all these tests. We conclude that emergent predictability is robustly supported by data.
@@
Furthermore, while we formulated our analysis in abstract terms such as the shape of a Pareto front, we can also confirm the result in more familiar terms by plotting the measured property against the prediction of the simplest models identified by our framework. Two examples are shown in Fig. 3F, G, which correspond to the points on the Pareto fronts labeled with arrowheads in Fig. 3A. At high richness, a simple two-parameter regression explains 85% of variance in butyrate production (Fig. 3G). In contrast, for low-richness communities, no model of comparable complexity can explain more than 49% of variance; the best-performing example is shown in Fig. 3F.
II.3 Mechanistic origin of emergent predictability
Why is this occurring? As emphasized above, our normalization procedure for the Pareto fronts ensures that the effect of Fig. 3 cannot be explained by simple self-averaging of the observable of interest.
@@ II.3 Mechanistic origin of emergent predictability
At least for the system of Clark et al., the emergent predictability appears to be related to the dominant structuring role of pH, which is one-dimensional. While the data is insufficient to demonstrate this directly (independent measurements of pH were conducted for only a subset of communities), we can obtain strong indirect evidence by leveraging the fact that four functional properties were measured simultaneously. These four measurements (the final concentrations of butyrate, acetate, succinate and lactate produced) define, for each community, a point in what we will call the “functional space”. Fig. 4A shows that at high community richness, these points are found along a lower-dimensional subspace. Indeed, for high-richness communities, 73% of variance is explained by PC1, compared to only 41% for low-richness communities (Fig. 4B). This dominant axis of variation is strongly correlated with community pH (Fig. 4C), consistent with Ref. [33] identifying pH-mediated interactions as a key factor structuring these communities. This observation explains the sudden drop in the Pareto fronts observed in Fig. 3A-D. Indeed, any combination of species whose abundance is correlated with pH will explain a substantial fraction of variance in any of the observables that have a loading on the dominant principal component of Fig. 4. As a result, even a little compositional information can be sufficient to improve predictive power, which is what the Pareto front indicates.
This argument reinforces the conclusion of Fig. 2 that the shape of the Pareto fronts observed in Fig. 3 is not due simply to the number of interacting species being large, but reflects the biology of the underlying communities. Crucially, the analysis of Fig. 3A-D used no knowledge of pH. Thus, our framework may provide a prior-knowledge-agnostic way for discovering the existence of strong ecosystem-structuring factors.
@@
For specifically Fig. 3A (butyrate), the results of Clark et al. [33] provide some additional mechanistic insight. Of the 25 strains in this study, five are capable of producing butyrate. However, in high-richness communities, butyrate production is dominated by one of them, Anaerostipes caccae (AC). And indeed, the first point on the high-richness Pareto front in Fig. 3A is the grouping {AC, all other strains}, and provides an excellent prediction for the amount of butyrate produced (Fig. 3G). In contrast, in low-richness mixtures, all five strains contribute to the production, and no single compositional feature is similarly predictive.
@@
Clark et al. hypothesized that this transition is associated with AC shifting at high community richness to an alternative metabolic pathway, utilizing lactate rather than sugars, a physiological switch mediated by sensing pH and/or the depletion of key resources. Identifying candidate mechanistic explanations for the other properties, and validating these hypotheses, would require new experiments. It is conceivable that the observations of Fig. 3A-D are all a consequence of specific physiological features of individual community members. However, the common outcome observed across two independent datasets suggests that emergent predictability as quantified here may be more general than a mechanistic idiosyncrasy of specific taxa.
Figure 4: Emergent simplicity in functional space. A: Principle component analysis of the four-dimensional functional measurements of Clark et al. (butyrate, acetate, lactate, succinate) indicates that high-richness samples form a lower-dimensional subspace. Principal axes computed for the dataset as a whole; mid-richness communities omitted from plotting for clarity; see SI for full dataset. B: PCA spectra computed separately for the two tiers of community richness. For high-richness samples, the dominant component explains 73% of variance, contrasted with 41% for low-richness. C: Coloring the PCA plot by measured pH reveals the dominant axis of functional variation is correlated with community pH. Axes same as in A, but including samples of all richness. In gray, communities for which pH was not measured.

### v3

@@ II Results
II.1 Data selection
@@
To apply our framework, we need a dataset satisfying several conditions. First, it must follow the combinatorial assembly protocol of Fig. 1A, with communities assembled from a fixed pool of strains. Surveying the literature, we identified 9 candidate datasets (across both microbial and non-microbial contexts; see Supp. Table S1). However, for our purposes it is essential that the assayed communities span a range of richness, so that at least two tiers of community diversity can be defined and compared. Additionally, if $N$ is the number of samples and $S$ the number of strains in the pool, $N/S$ should be sufficiently large (for each richness tier) to allow training the model from our chosen model class. Our choice of linear regressions as model class makes this last condition the most permissive and identifies two datasets whose structure allows our question to be asked. These are the data of Clark et al. [11], who assayed communities of 1-25 gut-derived taxa and measured the production of butyrate, acetate, lactate and succinate; and the work of Kehe et al. [33], who used a droplet-based platform to screen communities of 1-14 isolates, measuring the abundance of a model plant symbiont H. frisingense as the property of interest. For each dataset, we resolve three tiers of community richness: low richness, 151-5 strains; middle richness, 101510-15 strains; and high richness, 212521-25 strains. Together, these two datasets provide measurements of five properties we can examine for evidence of emergent simplicity. To construct the information-prediction Pareto fronts defined above, we use a Metropolis-Hastings algorithm to search the space of possible coarsenings (for details, see Methods).
II.2 Self-averaging is generic and ubiquitous
Figure 2: Self-averaging is generic. A: Coefficient of variation (CV) for each of the properties measured in Refs. [11] and [33], computed across communities of a given richness tier. The decreasing trend indicates self-averaging. Legend labels: Bu == Butyrate, Ac == Acetate, La == Lactate, Su == Succinate, FA == Focal species abundance. Shading marks $\pm 2$ SD using standard bootstrapping with 1000 sub-samplings. B: CV of synthetic properties generated from a null model where all species contribute proportional to their abundance with a weight drawn from a unit uniform distribution (see text). Three synthetic properties were computed using either shuffled abundance data from Refs. [11, 33] or by drawing random abundances from a Gaussian with mean and variance both set to 1. Null models reproduce the signature decrease of CV with richness, demonstrating that self-averaging is a generic expectation. Shading as in (A).
Figure 3: Coarse-grainability arises from community structure. A: Difference between the actual Pareto front and null ($\Delta$) as a function of description complexity, shown for empirical community properties measured in Refs. [11] and [33], and contrasted with model-derived properties (community invadability) in two standard ecological models: random-interaction CRM and GLV. $\Delta>0$ corresponds to coarse-grainability. See inset for the actual Pareto front and null comparison in the case of Butyrate as the property of interest. For CRM and GLV, the shading represents the quartiles in $\Delta$ across 10 random invaders. For empirical curves, error bar shading is omitted for clarity but is provided in the Supplementary Text (Fig. S8) Abbreviations as in Fig. 2. B: The simplest model capturing the fact that species traits are non-random: species’ contributions to a synthetic property (contribution weight) are assumed to be phylogenetically correlated, with conservation depth a tunable parameter (see Methods). The phylogenetic tree corresponds to the 25 species assayed by Clark et al. (reproduced from Ref. [11]). Leaf coloring illustrates the case where contribution weight is deeply conserved. C: $\Delta$ computed for a synthetic community property generated from the Clark et al. abundance data using phylogenetically correlated weights with conservation depth set to the genus or phylum level (see Methods). As expected, increasing conservation depth parameter leads to greater coarse-grainability. For comparison, the dashed line shows coarse-grainability for a synthetic property generated from central carbon metabolism gene profiles of the same 25 species predicted by PICRUSt2 (see Methods). Lines and shading represent median and quartiles across 10 replicate weight samplings.
For all five properties measured across both datasets, we see that their coefficient of variation decreases with community richness (Fig. 2A). This observation means their value is more reproducible across high-richness communities than it is among low-richness communities, and as such, could be described as a form of “emergent simplicity.” However, this same behavior is observed in the simplest of null models that retain no biological structure: self-averaging is preserved even if we randomly reshuffle the measured abundances, or use a purely synthetic property to which each species contributes independently (Fig. 2B; see Methods for details). This is not surprising, since both central-limit averaging and the increasing compositional similarity of high-richness samples (drawing $S>20$ species out of a pool of only 25) naturally contribute to this effect. Thus, when talking about emergent simplicity, it is essential that our terms are sufficiently well-defined to avoid any potential confusion with effects of trivial origin. To this end, from now on, we will ignore any absolute changes in the variance, assessing instead the fraction of variance explained by various models, and examining the shape and behavior of the Pareto front in these normalized units (Fig. 1E,F). Additionally, our null-model comparisons will employ the same sampling structure as the data, and whenever appropriate, the actual measured abundance tables, so that the compositional similarity of samples is fully retained.
II.3 Coarse-grainability arises from community structure
We say that a community property $Y$ is coarse-grainable if coarse compositional representations are surprisingly predictive of $Y$, compared to our null expectation. To establish the null expectation, we construct a synthetic property that uses the measured abundance table and preserves some statistics of the microscopic model fit to the actual property of interest, but assumes that species’ contributions are random and uncorrelated, eliminating any notion of biological similarity or relatedness between them (for details, see Methods). By construction, the Pareto front for this synthetic property has the same endpoints as the Pareto front of interest (Fig. 3A, inset). We quantify coarse-grainability by measuring the difference between the actual Pareto front and the null comparison, as a function of description complexity; positive values indicate that $Y$ is coarse-grainable. Applying this methodology to the two datasets, we find that all 5 observables across both datasets show a strong signature of coarse-grainability (Fig. 3A). Note that coarse-grainability is defined without reference to varying community richness; the analysis in Fig. 3A is for high-richness communities ($S>20$; for Kehe et al. $S>10$).
Coarse-grainability is clearly a useful feature for understanding complex ecosystems. It is therefore worth noting that some of the dominant theoretical frameworks of high-diversity ecology are missing this property. To illustrate this, we consider two classic frameworks widely used to model high-diversity ecosystems: the Generalized Lotka-Volterra model (GLV) with a random interaction matrix, and a Consumer-Resource model (CRM) with a random resource consumption matrix, choosing “community invadability” as the property of interest (specifically, $Y$ is defined as the post-invasion abundance of some specified strain, initially absent from the community). This choice is convenient because it is a relevant and non-trivial community property that can be accessed in both modeling frameworks. The model parameters and simulation protocol are described in the Methods, but the outcome of the analysis is that neither model exhibits coarse-grainability (Fig. 3A, shaded).
Figure 4: Both empirical datasets exhibit emergent predictability. A: The Pareto front in error-complexity plane, computed for the butyrate measurements of Clark et al. [11]. Lines show the fraction of variance left unexplained (FVU) by the best model within a given description complexity, when dataset is restricted to communities of low richness ($S\leq 5$, orange), medium richness (gray) or high richness ($S>20$, purple). Points indicate the median out-of-sample FVU computed over 100 random 50-50 splits of the data into training and testing sets; error bars are the 1st and 3rd quartiles. The shift of the Pareto front indicates that at higher richness, coarse descriptions become more predictive (cf. Fig. 1F). Panels B-C illustrate this result by plotting the predicted vs. measured concentration for the two models indicated by the arrowheads. At high richness, a highly coarsened model explains 85% of variance in butyrate production (C). In contrast, for low-richness communities, even the most predictive model of comparable complexity explains only 49% of variance (B). D-G: Same as (A) for the four other properties measured in Refs. [11] and [33]. The Kehe et al. dataset (panel G) only assayed 14 isolates and is missing the high-richness tier ($S>20$ strains). H: To quantify emergent predictability (EP), we measure the signed area $\mathcal{A}$ between the low-richness and highest-available-richness Pareto fronts of previous panels (highlighted in G). The quantification confirms that all measured properties exhibit emergent predictability. This is in contrast to all theoretical models we considered, including those that exhibit coarse-grainability (CRM, random-interaction consumer-resource model; GLV, random-interaction Lotka-Volterra; P2, PICRUSt2 central carbon metabolism; Phy, phylogenetically conserved contributions; other abbreviations as in Fig. 2). Error bars show quartiles of the respective distributions. For empirical data, they represent the uncertainty of Pareto front estimation; for models, they include variability across random model instances (see Methods).
The origin of this contrast is easy to pinpoint: real species and interactions are not random. If relatedness of species is taken into account, then even the simplest models can exhibit coarse-grainability. To illustrate this, we construct a simple model in which the only assumption is that closely related species contribute similarly to a community-level property. Using the abundance data from Clark et al., we define a synthetic community property to which each species contributes independently, but with weights that are correlated according to phylogenetic relatedness (Fig. 3B shows an example). A single tunable parameter controls the phylogenetic scale at which trait correlations decay, capturing the depth of conservation (see Methods). As expected, increasing the depth of trait conservation leads to greater coarse-grainability (Fig. 3C). Of course, the traits relevant to the measurements of Clark et al. and Kehe et al. are unlikely to have a single relevant scale of conservation depth. As an example of more realistic structure, we also include a coarse-grainability curve for a different synthetic property, constructed from the presence of carbon metabolism genes (KEGG orthologs) predicted by PICRUSt2 [17] for the same set of species; see Methods for details. The shape of this curve confirms that a realistic amount of structure is sufficient to generate coarse-grainability. We note that the scale of the $y$ axis in Fig. 3C depends on arbitrary modeling choices and is not meant to be directly compared to Fig. 3A; our focus is on the coarse-grainability $\Delta$ becoming positive at intermediate complexity.
This analysis demonstrates that even minimal models incorporating phylogenetic signal can exhibit coarse-grainability. Some theoretical frameworks intentionally omit this relatedness structure, and thus miss this effect, but in any empirical scenario (or in models using realistic trait matrices, e.g., CRM based on measured resource preferences [28, 24]), at least some degree of coarse-grainability is to be expected.
II.4 Emergent predictability is surprising and demands explanation
Finally, we turn to emergent predictability, encoded in the behavior of the Pareto front under changes of richness. For each dataset, we resolve three tiers of community richness. We apply the same search algorithm as above to infer the respective Pareto front of each richness tier. The results are shown in Fig. 4A,D-G.
@@ II.4 Emergent predictability is surprising and demands explanation
Remarkably, we find that all five properties exhibit emergent predictability. We emphasize that the Pareto front is not an abstract concept: each point corresponds to an explicit model, and we consistently find that coarse models become increasingly predictive in communities of higher richness. As an illustration, Fig. 4B&C plot the measured butyrate concentration against the prediction of the simplest models identified by our framework, labeled with arrowheads in A. At high richness, a coarsening exists whereby a simple two-parameter regression explains 85% of variance in butyrate production (Fig. 4C). In contrast, for low-richness communities, no model of comparable complexity can explain more than 49% of variance; the best-performing example is shown in Fig. 4B.
As a control, this behavior is not observed if the values of $Y$ or species abundances are randomly permuted (see Supplementary Text, Figs. S6S7), indicating that the behavior observed in Fig. 4 is not an artifact of our analysis. The additional controls we performed include subsampling the data, perturbing the groupings forming the putative Pareto front, and considering alternative metrics of complexity $I(\Psi)$; see Supplementary Text, Figs. S3S4S1, respectively. We conclude that emergent predictability is robustly supported by data.
Crucially, however, none of the standard models show this behavior. To see this, we quantify emergent predictability by measuring the area contained between the high-richness and low-richness Pareto fronts (see Methods). Fig. 4H summarizes the contrast between the empirical results and all the theoretical models discussed so far. Each of these models exhibits self-averaging, and multiple models exhibit some degree of coarse-grainability, provided they retain some notion of relatedness between species (PICRUSt; phylogenetically conserved contributions). However, none of these models exhibit emergent predictability. (See also Supplementary Text section C.1.)
II.5 Mechanistic origin of emergent predictability
@@
At least for the system of Clark et al., the emergent predictability appears to be related to the dominant structuring role of pH, which is one-dimensional. While the data is insufficient to demonstrate this directly (independent measurements of pH were conducted for only a subset of communities), we can obtain strong indirect evidence by leveraging the fact that four functional properties were measured simultaneously. These four measurements (the final concentrations of butyrate, acetate, succinate and lactate produced) define, for each community, a point in what we will call the “functional space”. Fig. 5A shows that at high community richness, these points are found along a lower-dimensional subspace. Indeed, for high-richness communities, 73% of variance is explained by PC1, compared to only 41% for low-richness communities (Fig. S11). This dominant axis of variation is strongly correlated with community pH (Fig. 5A, inset), consistent with Ref. [11] identifying pH-mediated interactions as a key factor structuring these communities. This observation likely explains the sudden drop in the Pareto fronts observed in Fig. 4A,D-G. Indeed, any combination of species whose abundance is correlated with pH will explain a substantial fraction of variance in any of the community properties that have a loading on the dominant principal component of Fig. 5. As a result, even a little compositional information can be sufficient to improve predictive power, which is what the Pareto front indicates. Crucially, the analysis of Fig. 4 used no knowledge of pH. Thus, our framework may provide a prior-knowledge-agnostic way for discovering the existence of strong ecosystem-structuring factors.
Figure 5: Mechanistic origin of emergent predictability. A: Principle component analysis (PCA) of the four-dimensional functional measurements of Clark et al. (butyrate, acetate, lactate, succinate) indicates that high-richness samples form a lower-dimensional subspace. Coloring the PCA plot by measured pH (inset) reveals that the dominant axis of functional variation is correlated with community pH. Principal axes computed for the dataset as a whole; mid-richness communities are omitted for clarity (but included in the inset). B: A minimal model that can reproduce emergent predictability. The key ingredient is a feedback loop (of strength $\theta$) with a shared environmental variable pp. Species differ by their preferred interval of $p$ which allows them to be active; see text for details. Species’ contribution strengths to the property of interest $Y$ are random and uncorrelated (no special structure is assumed). C: The model exhibits emergent predictability; shown is an example for feedback strength $\theta=0.8$. Richness tiers match Clark et al., with 800 simulated communities per tier (see Methods). D: In the model, emergent predictability is controlled by feedback strength $\theta$. As in Fig. 4, emergent predictability is quantified as the signed area between low- & high-richness Pareto fronts (shaded in panel C). Error bar shading shows quartiles over 10 random model instances. The feedback strength corresponding to the example shown in panels C & E is highlighted. E: Scatter plot of the community property $Y$ versus the environmental variable $p$ for the same simulated dataset as used in panel C; colors as in (A). The plot offers an intuitive explanation of the origin of emergent predictability in the model. At higher richness (purple), the point cloud is tighter, but the tightening is anisotropic, allowing $p$ (and, by extension, $p$-dependent compositional features) to become more predictive of $Y$. F: A cartoon graphical summary of the proposed general mechanism illustrated by our model. Communities are represented as points in an abstract “functional space.” As richness increases, communities generally become more functionally similar due to generic forms of self-averaging. However, if physiological or environmental feedbacks oppose self-averaging along some axes of community variation (e.g., the $p$ axis in our model), then at high richness, such axes will become responsible for an increasingly larger fraction of variance in related properties.
@@
For specifically Fig. 4A (butyrate), the results of Clark et al. [11] provide some additional mechanistic insight. Of the 25 strains in this study, five are capable of producing butyrate. However, in high-richness communities, butyrate production is dominated by one of them, Anaerostipes caccae (AC). And indeed, the first point on the high-richness Pareto front in Fig. 4A is the grouping {AC, all other strains}, and provides an excellent prediction for the amount of butyrate produced (Fig. 4C). In contrast, in low-richness mixtures, all five strains contribute to the production, and no single compositional feature is similarly predictive.
@@
Clark et al. hypothesized that this transition is associated with AC shifting at high community richness to an alternative metabolic pathway, utilizing lactate rather than sugars, a physiological switch mediated by sensing pH and/or the depletion of key resources. Identifying candidate mechanistic explanations for the other properties, and validating these hypotheses, would require new experiments. It is conceivable that the observations of Fig. 4A, D-G are all a consequence of some idiosyncrasies of specific taxa. However, the common outcome observed across two independent datasets suggests that emergent predictability as quantified here may have a more general origin.
II.6 Candidate general mechanism: ecological feedback that opposes self-averaging
We will now describe and support a candidate general mechanism for emergent predictability to arise. Somewhat counter-intuitively, we propose that emergent predictability arises when some physiological or environmental feedback opposes statistical self-averaging along certain axes of community variation. To illustrate, we construct a simple model (Fig. 5B) inspired by Refs. [59, 2, 50]. Let species $i$ be characterized by metabolic activity $a_{i}$. The metabolic activity defines its growth, scales its contribution to a collective property of interest $Y$, but also its influence on a shared environmental variable $p$ (e.g., pH, osmolarity, oxygen content, toxin concentration):
$\dot{n}_{i}$ $=n_{i}a_{i}$ population growth
$\dot{r}$ $=-\sum_{i}a_{i}n_{i}$ nutrient depletion
$\dot{p}$ $=\sum_{i}a_{i}n_{i}c_{i}$ dynamics of $p$
$\dot{Y}$ $=\sum_{i}a_{i}n_{i}y_{i}$ $\text{dynamics of $Y$}.$
We will assume no special structure in species’ contributions to either $p$ or $Y$: both $c_{i}$ and $y_{i}$ will be drawn randomly and independently from Gaussian distributions. However, crucially, we will assume that the activities $a_{i}$ are impacted by feedback from $p$. Specifically, each species $i$ has a different interval of $p$ that it prefers, and can only be active within a finite range of a preferred value (denoted $\rho_{i}$):
$a_{i}=r~F(|p-\rho_{i}|).$
Here, $r$ is nutrient availability, and $F$ has a finite width $\simeq\Delta p$ (we pick $F$ to be a truncated Gaussian; see Methods). As an example, a species acidifying its own environment can poison its own growth in monoculture, even if nutrient is still available [58]. The parameter $\theta=1/\Delta p$ encodes the feedback strength: large $\theta$ corresponds to species highly sensitive to changes in $p$ (a narrow activity range), whereas $\theta=0$ corresponds to $p$ being irrelevant. This feedback, assumed to be mediated by species physiology, is the key ingredient of our model. For more details, see Methods.
Fig. 5C confirms that this model displays emergent predictability (simulation parameters are listed in Methods). As we expected, the magnitude of the effect is controlled by feedback strength $\theta$ (Fig. 5D). Intuitively, this can be understood as follows. In the analytically solvable limit $\theta\rightarrow 0$ (no feedback), it is easy to see that $p$ is irrelevant, the species’ contributions to $Y$ are subject to simple self-averaging, and there is no emergent predictability (for more discussion, see Supplementary Text, section C.1). At non-zero $\theta$, self-averaging continues to play a role leveling out the quirks of individual species within a tier of similar $p$ preference. However, along the $p$ axis, the $p$-mediated ecological feedback opposes self-averaging. For instance, a species that is strong driver of $p$ (large $c_{i}$) will always drive this environmental variable to the upper limit of its viable range, which is a fixed value set by species physiology ($\simeq\rho_{i}+\Delta p$) and does not depend on richness. As a result, the axis of variability associated with $p$ preference maintains its relevance while other differences self-average, and so becomes responsible for an increasingly larger fraction of the total variance of the property of interest (illustrated in Fig. 5E, F). A more detailed analysis of this mechanism can be found in the Supplementary Text.
The feedback of the kind we described is expected to be widespread [50], and Fig. 5D demonstrates it is sufficient to endow a system with emergent predictability. Thus, our analysis suggests that emergent predictability is a nontrivial and exciting feature that might well be ubiquitous in natural communities.